OpenStudy (anonymous):

I dont get what its asking can someone help. Solve P=4(3a-b) for a in terms of P and b.

OpenStudy (anonymous):

cat fight?

OpenStudy (anonymous):

if a(b)=c Then c/b=a ....

OpenStudy (anonymous):

Basically instead of expressing P=f(a,b) in terms of a and b, you have to express a=f'(P,b) in terms of P and b

OpenStudy (anonymous):

still confused -_-

OpenStudy (anonymous):

P=4(3a-b) or P/4=3a-b or (P/4) +b=3a so a=P/12 +b/3

OpenStudy (anonymous):

it wants you to put p and b on the same side. generalize and use algebra

OpenStudy (merchandize):

multiply out. P = 12a - 4b Now move -4b to the left hand side because you want to isolate 12a +4b +4b So you add 4b to both sides P+4b = 12a Divide by 12 on both sides to get a P+4b/12 = a Reduce 4b/12 by dividing by 4 P+ b/3 = a Final answer

OpenStudy (anonymous):

merchandize, the first term would be P/12 after dividing by 12

zepdrix (zepdrix):

Yah the wording was a little confusing Solo, they're just telling you to "Solve for a" :)

OpenStudy (merchandize):

okay

OpenStudy (merchandize):

the question says that calculate a in terms of p and b.......so we just calculate the value of a in terms of p and b....

OpenStudy (merchandize):

the value of a will be only the answer